ON LOCAL SOLUTIONS TO A RADIAL p-ELLIPTIC EQUATION ON (0, ∞)

KOICHI UCHIYMA

Abstract : Reviewing power series description of analytic singularities of local solutions to a radial p-Laplace equation near a point σ in (0, ∞), I will discuss power series construction of solutions U(r) near the ends of (0, ∞) to a radial p-elliptic equation (r^{n−1}a(r)^{p/2}|U_{r}|^{p−2}U_{r})_{r} + λr^{n−1}ρ(r)|U|^{q−2}U = 0 for 1 < p, q < ∞ and nonzero real λ. A Briot-Bouquet type theorem of two varibales is used for these purposes.